4 Assume that a purely competitive firm has the schedule of the average and marginal costs given in the table below

-------------------------------------------------------

OUTPUT AFC AVC ATC MC

--------------------------------------------------------

1          $300 $100 $400 $100

2           150    75    225    50

3           100    70   170     60

4            75     73   148     80

5            60     80   140   110

6            50     90   140   140

7            43    103   146 180

8            38    119   156 230

---------------------------------------------------------

a. At a price of $68, the firm will produce _____ units of output. The firm's economic profit is ________.

b. At a price of $80, the firm will produce ______ units of output. The firm's economic profit is ______.

Will the firm break-even at this price? _____. If not, what will be this firm’s break-even price? ______.

Explain why. ______________________________________________________________________

c. At a price of $190, the firm will produce ________ units of output. Will $190 be the long run price for

this firm? _____ Explain your answer. _________________________________________________

If this is not the long run price, what will be the long run price? _____.
Answer #1

(a) Profit maximizing condition is when P=MC.

At a price of $68, the firm will produce 3 units of output because P>MC at Q=3 and P<MC at Q=4. The firm's economic profit is (68)(3)-(170)(3)= 204- 510= -$306 (i.e economic loss).

(b) At a price of $80 , the firm will produce 4 units of output (because P=MC at Q=4). The firm's economic profit is (80)(4)-(148)(4)= 320-592= -$272 (i.e economic loss).

No, the firm will not break-even at this price because at break-even point ATC=MC. Therefore, the firm will break-even at the price of $140.

(c) At a price of $190 , the firm will produce 7 units of output (because P>MC at Q=7 and P<MC at Q=8).

No, the firm will not break-even at this price because at break-even point ATC=MC. Therefore, the firm will break-even at the price of $140.